=== 2019/2/4(Mon.) ===

=== Weekend ===

=== Schedule ===
  * Current status as of 1/22: IR and Green reached X end PD.
  * --(1/22: Surface inspection for PR mirrors.)-- Cleaning will be done after 1/31.
  * --(1/23: Surface inspection for SR mirrors.)--
  * --(1/24-25: Cleaning HR of SR3, SRM.)--

  * --(1/23: BS is aligned to Y-end PD by IR.)--
  * --(1/27: ITMY is aligned to REFL by IR.)--
  * --(1/27: ITMX is aligned to REFL by IR.)--
  * --(1/27: Beam position of IR at SR3 is confirmed how far from the center.)--

  * --(1/28-29: SR3 is aligned to SR2 center by IR.)--
  * --(1/28: Alignment of SR2 -> SRM by IR.)--
  * --(1/29: Green beam is injected from SR2 AR surface to Y end PD.)--

  * --(1/31: Removing bellows at PR and SR area.)--
  * --(1/31-2/1: Cleaning HR of PR3 and HR, AR of SR2)--

  * 2/4-8: SRM -> OMMT by IR -> OMMT -> OSTM by IR, OSTM -> OMC by IR
  * 2/4-8: Alignment of ETMY -> POS by Green

  * 2/12-15: Closing door for central area, pumping down
   * needs drift monitor for VIS

==== Note ====

 * Report klog, and gokagra ML if urgent things.

==== Report of work ====
 *MIR:
 * SRM: It seems to be that mirror rotated 180degree.
  * It is still not perfectly sure actually SRM was rotated, it could be design itself or mistake of position of something downstream.
  * In any case, we need to correct something.
  * Type B team will look at probably today and tomorrow.
 * COMM: Needs some alignment downstream SRM due to the mystery above.
  * IR did not hit SR2 center. It will be checked.

 * IOO: For high power test, we could inject less 10W due to saturation of FSS.
 * MIR: HR surface of EXMX has some dusts. We should discuss whether we should clean or we should leave it with some information of other test masses.



==== Work plan ====